& Solve the IVP: X=AX where - A=(-3 5 7 and X(0)=(=;)
Answer #1

2 2 sole Ivp X-Ax where AP5]. (c) (6) = first we find egenvalues IA-da 120 13:15) -0 (3-) (1-x) (5) (-1) = 0 -3-x + 3x + x² +This, the solution that we get from first eigenvalue and first eigen vector is xict) eltist (2-2) =ete it (2-2) etcostuusint>20-2=-1 2(-1)-2=-1 = -2- =-1 -2+1=ca - ca- 0,= -1,2=-) -gensol .- (t): et pole sint ) et les con : 1 - 20 test - etsint-net s